# MiSS: Revisiting the Trade-off in LoRA with an Efficient Shard-Sharing Structure

### v11

@@ 5 Experiments
In this section, we evaluate the performance of MiSS on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset . Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM.
@@
Method Trainable MNLI SST-2 CoLA QNLI MRPC
LoRA 0.236% 85.63±0.01 94.03±0.02 62.40±0.71 91.37±0.97 87.98±0.23
PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24
MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
@@
Model Strategy Trainable GSM8K Math HumanEval Mbpp
Llama2-7B (Touvron et al., 2023) LoRA 89.9M 40.75 5.22 17.74 35.15
DoRA 91.3M 42.93 6.51 21.95 36.53
PiSSA 89.9M 43.89 6.92 22.15 37.84
MiSS 87.0M 48.16 8.58 23.63 36.81
RWKV 6-7B (Peng et al., 2024) LoRA 88.1M 38.13 6.06 - -
PiSSA 88.1M 40.48 6.12 - -
MiSS 88.1M 41.73 6.52 - -
Mistral-7B (Jiang et al., 2023) LoRA 94.4M 62.85 15.82 35.71 46.11
DoRA 95.8M 63.68 13.60 38.41 48.73
PiSSA 94.4M 67.01 18.13 41.28 51.37
MiSS 87.0M 68.92 18.85 42.07 61.33
Llama2-13B (Touvron et al., 2023) LoRA 250M 56.18 12.60 31.79 37.82
DoRA 252M 61.56 13.60 33.50 39.25
PiSSA 250M 66.64 13.82 33.57 46.03
MiSS 255M 68.64 15.74 38.15 47.91
Qwen3-4B (Yang et al., 2025) LoRA 74.3M 84.38 15.20 73.27 78.32
DoRA 75.4M 85.11 21.73 74.20 78.77
PiSSA 74.3M 85.78 26.00 75.01 78.04
MiSS 70.1M 85.52 34.82 74.48 78.05
5.1 Experiments on Natural Language Understanding
5.2 Experiment on Natural Language Generation
Setup.
To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval (Chen et al., 2021) and Mbpp (Austin et al., 2021). The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
Results.
(a) Loss-Token
(b) Loss-Time
The results (See Table  3) show that MiSS consistently achieves the best or near-best performance across different model scales and diverse tasks. For example, on Llama2-7B, MiSS significantly surpasses LoRA on GSM8K (48.16 vs. 40.75) and also improves HumanEval and Mbpp with scores of 8.58 and 36.81, respectively. As the model scales up to Llama2-13B, MiSS maintains its advantage, particularly on code generation benchmarks such as HumanEval and Mbpp, outperforming both PiSSA and DoRA. More importantly, on the latest large-scale model Qwen3-4B, which is trained on richer data, MiSS achieves the highest Math score of 34.82, far exceeding other methods, while remaining competitive on GSM8K and Mbpp. These findings demonstrate that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
@@ 5.3 Effect of Rank $r$
Table 5: Time and space complexity. e denotes efficient form.
Method Space Time
Full $O(dk)$ $O(bld(d+k))$
LoRA $O(dr+rk)$ $O(blr(d+k))$
MiSS $O(dr)$ $O(bldk)$
MiSSe $O(dr)$ $O(blr(d+\frac{k}{r}))$
5.4 Pareto plot
The emergence of PEFT techniques is motivated by dual objectives: mitigating GPU memory constraints and exploring more efficient model architectures. Nevertheless, numerous contemporary studies disproportionately focus on ultimate performance benchmarks, overlooking critical practical considerations like computational efficiency and training duration—an emphasis that clearly diverges from the original rationale for PEFT. In this section, we undertake a multi-dimensional investigation into the relationships among computational overhead, efficiency, and performance for diverse models. Leveraging the official Hugging Face PEFT (Mangrulkar et al., 2022) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
@@ 7 Limitations and Future Work
Through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS. The strong performance achieved on large language models gives us substantial confidence in the potential of this approach. We believe that MiSS can be readily applied to a wide range of multimodal tasks.
@@ Appendix A Appendix
A.1 RWKV7
@@
Appendix B Settings of Experiments in No Free Lunch

### v12

@@ 5 Experiments
In this section, we conduct a comprehensive set of experiments to validate the effectiveness and generalizability of MiSS across diverse domains. We assess performance on a wide range of tasks, including language, image, and video benchmarks. Specifically, we evaluate Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset, and Natural Language Generation (NLG) capabilities by fine-tuning various large language models (LLMs). We extend our evaluation to multimodal settings using the VTAB-1K benchmark to demonstrate the robust adaptability of MiSS beyond textual domains. Furthermore, we provide a detailed analysis of the Pareto frontier (Section 5.3) to definitively illustrate MiSS’s superior computational efficiency and minimal hardware overhead when compared to existing Parameter-Efficient Fine-Tuning (PEFT) methods.
5.1 Superior Performance across Language and Vision Domains
MiSS demonstrates exceptional versatility, maintaining a commanding lead or highly competitive performance across diverse benchmarks in both the language and vision domains. (Setup B)
Natural Language Understanding (NLU).
On the GLUE benchmark (Table 4), fine-tuning RoBERTa-base with MiSS showcases notable strength. It achieves an outstanding result on the challenging CoLA dataset (72.86), significantly surpassing LoRA and PiSSA. This performance indicates superior data-fitting capabilities and faster convergence on complex linguistic tasks.
@@ 5.1 Superior Performance across Language and Vision Domains
Method Trainable MNLI SST-2 CoLA QNLI MRPC Avg
LoRA 0.236% 85.63±0.01 94.03±0.02 62.40±0.71 91.37±0.97 87.98±0.23 84.28
PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24 85.23
MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60 86.35
Natural Language Generation (NLG).
Across five mainstream LLMs (Llama2, Mistral, RWKV, Qwen3), MiSS consistently achieves the best or near-best average performance (Table 5). Notably, it demonstrates substantial gains in complex reasoning tasks, recording the highest Math score (34.82) on Qwen3-4B and the highest average score (47.79) on Mistral-7B. These findings highlight that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
@@
Model Strategy Trainable GSM8K Math HumanEval Mbpp Avg
Llama2-7B (touvron2023llama) LoRA 89.9M 40.75 5.22 17.74 35.15 24.72
DoRA 91.3M 42.93 6.51 21.95 36.53 26.48
PiSSA 89.9M 43.89 6.92 22.15 37.84 27.70
MiSS 87.0M 48.16 8.58 23.63 36.81 29.30
RWKV 6-7B (peng2024eagle) LoRA 88.1M 38.13 6.06 - - 22.10
PiSSA 88.1M 40.48 6.12 - - 23.30
MiSS 88.1M 41.73 6.52 - - 24.13
Mistral-7B (jiang2023mistral7b) LoRA 94.4M 62.85 15.82 35.71 46.11 40.12
DoRA 95.8M 63.68 13.60 38.41 48.73 41.10
PiSSA 94.4M 67.01 18.13 41.28 51.37 44.45
MiSS 87.0M 68.92 18.85 42.07 61.33 47.79
Llama2-13B (touvron2023llama) LoRA 250M 56.18 12.60 31.79 37.82 34.60
DoRA 252M 61.56 13.60 33.50 39.25 36.98
PiSSA 250M 66.64 13.82 33.57 46.03 39.52
MiSS 255M 68.64 15.74 38.15 47.91 42.11
Qwen3-4B (yang2025qwen3technicalreport) LoRA 74.3M 84.38 15.20 73.27 78.32 62.79
DoRA 75.4M 85.11 21.73 74.20 78.77 64.95
PiSSA 74.3M 85.78 26.00 75.01 78.04 66.21
MiSS 70.1M 85.52 34.82 74.48 78.05 68.22
Vision Task
To validate the ability of MiSS to adapt to non-textual tasks, we conducted experiments on the VTAB-1K image and video benchmarks (Table 6). MiSS achieved an average accuracy of 88.02 on image tasks and 72.96 on video tasks, making it highly competitive with top-performing baseline methods like LoRA and DoRA. Crucially, this competitive performance is delivered with a significantly lower parameter budget ($\approx 0.4$ #TPs) compared to LoRA/DoRA ($\approx 0.8$ #TPs), confirming that the efficiency of MiSS transcends the language domain and is applicable to multimodal foundation models.
Table 6: Performance comparison on VTAB-1K image and video benchmarks.Results are adopted from SliceFine (kowsher2025slicefineuniversalwinningslicehypothesis).
Method Image Video
Caltech Flowers Pets Camel. Euro. Retino. KITTI Avg #TPs UCF101 Kinetics HMDB Avg #TPs
Full 89.92 97.41 85.87 81.65 88.12 73.62 77.93 84.93 85.83 92.30 55.23 65.79 74.99 86.65
VeRA 91.53 99.19 91.04 86.45 92.97 74.25 77.92 87.62 0.240 92.28 57.21 66.77 72.09 0.242
LoRA 92.03 99.18 90.92 87.73 92.65 74.23 80.42 88.08 0.833 93.88 57.81 67.37 73.02 0.835
DoRA 91.86 99.27 91.08 85.88 91.42 75.28 80.46 87.89 0.834 92.84 57.77 67.33 72.65 0.836
MiSS 92.14 99.23 91.05 86.28 92.83 73.71 80.91 88.02 0.414 93.82 57.75 67.31 72.96 0.415
@@ 5.2 Effect of Rank $r$
5.3 MiSS’s Superior Balance on the Pareto Frontier: Optimally Trading Off Efficiency and Performance
The emergence of PEFT techniques is motivated by dual objectives: mitigating GPU memory constraints and exploring more efficient model architectures. Nevertheless, numerous contemporary studies disproportionately focus on ultimate performance benchmarks, overlooking critical practical considerations like computational efficiency and training duration—an emphasis that clearly diverges from the original rationale for PEFT. In this section, we undertake a multi-dimensional investigation into the relationships among computational overhead, efficiency, and performance for diverse models. Leveraging the official Hugging Face PEFT (peft) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
@@ Appendix A Appendix
A.1 Additional Experiments
(a) Loss-Token
(b) Loss-Time
A.2 RWKV7
@@
Appendix B Settings of Experiments
NLU
NLG
To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA (yu2023metamath), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (cobbe2021gsm8k) and MATH  (yu2023metamath), with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback (zheng2024opencodeinterpreter), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval (chen2021evaluating) and Mbpp (austin2021programsynthesislargelanguage). The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
Vision Task
on VTAB-1K image classification using ViT-Base-Patch16-224
Appendix C Settings of Experiments in No Free Lunch